Part A and B
Part A Constants In a World Cup soccer match, Juan is running due north toward the goal with a speed of 740 m/s What is the magnitude of the balls velocity relative to Juan? relative to the ground. A teammate passes the ball to him. The ball has a speed of 12.2 m/s and is moving in a direction of 32.0 east of north, relative to the ground. UR m/s u a Previous Answers Request Answer X Incorrect; Try Again; 5 attempts remaining Part B What is the direction of the balls velocity relative to Juan? ° north of east
Answer #1

here,

the velocity of Juan , v1 = 7.4 m/s j

the velocity of ball , v2 = 12.2 m/s * ( sin(32) i + cos(32) j)

v2 = 6.47 i m/s + 10.35 j m/s

the velocity of ball relative to Juan , v21 = v2 - v1

v21 = 6.47 i m/s + 2.95 j m/s

a)

the magnitude of velocity of ball relative to Juan , |v21| = sqrt(6.47^2 + 2.95^2) = 7.11 m/s

b)

the direction , theta = arctan(2.95/6.47) = 24.5 degree North of East